OpenStudy (anonymous):

Can someone please explain this to me? Identify the maximum and minimum values of the function y = 3 cos x in the interval [-2π, 2π]. Use your understanding of transformations, not your graphing calculator.

OpenStudy (anonymous):

Do you know the minimum and maximum values of \(y=\cos x\) over the interval \([-2\pi,2\pi]\)?

OpenStudy (anonymous):

Honestly, I don't even understand what it's asking for. What does it mean by the interval [-2pi, 2pi]?

OpenStudy (anonymous):

\(-2\pi \leq x \leq 2\pi\) i.e. \(x\) is between \(-2\pi\) and \(2\pi\)

OpenStudy (anonymous):

Okay, that means on a graph. Do the signs always go that way? \[--> \le \]

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

you could write it \[2\pi \geq x \geq -2\pi\] as long as it means "in between"

OpenStudy (anonymous):

Oh, okay. I gotcha. So how do find what you said? y= cos x?

OpenStudy (anonymous):

Well, you could try putting in values of x or drawing it or, the most general method of finding minima or maxima of a function \(f(x)\) is to differentiate and set to zero \[\frac{\text{d}f}{\text{d}x} = 0\] so solving \(-\sin x = 0\) will give you values for \(x\) for which \(\cos x\) is maximum

OpenStudy (mathmale):

BG: I think you'd find it very helpful to actually graph y = 3 cos x on the given interval. From the graph it'd be immediately evident that this function has more than one max and more than one min within that interval. You cool with graphing the cosine function y=cos x? y=3cos x?

OpenStudy (anonymous):

Hold on a minute, MM. I will do that afterwards. I have to explain how I found my answer without the use of a calculator, but I'd love the tips on how to do this more easily afterwards! :)

OpenStudy (mathmale):

Broken Symm's most recent statement is helpful and accurate. But I'd still suggest graphing the function y=3cos x on [-2pi,2pi].

OpenStudy (anonymous):

What is d in that equation?

OpenStudy (anonymous):

*maximum or minimum I should have said

OpenStudy (mathmale):

I'm going to graph y=cos x on [0,2p] quickly. We really need to know by heart the shapes of the graphs of sin x, cos x and tan x, so that no calculator is needed to graph them.

OpenStudy (anonymous):

oh, \[\frac{\text{d}}{\text{d}x}\] is an operator Have you done any calculus?

OpenStudy (anonymous):

Nope. That's the class I'm in. Haha.

OpenStudy (anonymous):

Was that d/dx?

OpenStudy (mathmale):